Cushing Limited is considering the purchase of a new grading machine to replace the existing one. The existing machine was purchased three years ago at an installed cost of $20,000. It was being depreciated using the prime cost method with an effective life of five years. The existing machine is expected to have a usable life of at least five more years. The new machine costs $35,000 and requires $5,000 in installation costs. It will be depreciated using the prime cost method over five years. The existing machine can currently be sold for $25,000 without incurring any removal or clean-up costs. The firm pays 40% taxes on both ordinary income and capital gains. Calculate the initial investment associated with the proposed purchase of a new grading machine.
Answer #1

The initial investment will have following components:

  1. Purchase cost of new machine = $ 35,000
  2. Installation cost of new machine = $ 5,000
  3. Post tax salvage value on sale of old / existing machine:
    1. Pre tax sale value = $ 25,000
    2. Taxable basis =
      1. Purchase cost = 20,000
      2. Effective Life assumed = 5 years
      3. Annual depreciation under prime cost method = 20,000 / 5 = 4,000
      4. Accumulated depreciation in 3 years = 4,000 x 3 = 12,000
      5. Taxable basis = Purchase cost - accumulated depreciation = 20,000 - 12,000 = 8,000
    3. Gain on sale = Pre tax salvage value - taxable basis = 20,000 - 8,000 = 12,000
    4. Tax on gain = Tax rate x Gain = 40% x 12,000 = 4,800
    5. Post tax salvage value = pre tax salvage value - tax on gain on sale = 20,000 - 4,800 = 15,200

Thus, initial investment = Purchase cost of new machine + Installation cost of new machine - Post tax salvage value on sale of old / existing machine = $ 35,000 + 5,000 - 15,200 = $ 24,200
: